please help.
suppose a bag contains 6 marbles.one is white,two are red,and three are yellow.each time a marble is taken out,it is replaced.find P (red and yellow)


thanks in advance for the help.

Answer by stanbon(75887)

suppose a bag contains 6 marbles.one is white,two are red,and three are yellow.each time a marble is taken out,it is replaced.find P (red and yellow)
---------------
Since the selections are replaced the results are independent.
--------------
P(red AND yellow) = P(red)*P(yellow) = 2/6 * 3/6 = 6/36 = 1/6
==================
Cheers,
Stan H.